OpenStudy (anonymous):

ILL FAN AND GIVE A MEDAL TO WHOEVER CAN GIVE ME THE CORRECT ANSWER AND HELP WITH THIS!:) Find the exact value for sin(x + y) if sin(x) = -4/5 and cos(y) = 15/17. Angles x and y are in the fourth quadrant. Thanks in advance to whoever does!

OpenStudy (kc_kennylau):

remember \(\sin^2x+\cos^2x=1\) :D (corollary: \(\sin x=\sqrt{1-\cos^2x}\) and \(\cos x=\sqrt{1-\sin^2x}\) and you can use the formula \(\sin(x+y)=\sin x\cos y+\cos x\sin y\)

OpenStudy (kc_kennylau):

hope that helps you :D

OpenStudy (anonymous):

omgosh, i dont know what any of that means :c lol I'm great in pretty much every other subject, but trigonometry will be the death of me. haha Thanks for trying though !

OpenStudy (kc_kennylau):

just use the formula \(\sin(x+y)=\sin x\cos y+\cos x\sin y\) :)

OpenStudy (kc_kennylau):

and substitute the values into that formula

OpenStudy (anonymous):

so sin(-4/5+15/7)=my answer??

OpenStudy (kc_kennylau):

no :/

OpenStudy (kc_kennylau):

\(\sin(x+y)=\sin x\cos y+\cos x\sin y=-\frac45*\frac{15}{17}+\cos x\sin y\)

OpenStudy (anonymous):

that's the equation? how in the world did you get that??

OpenStudy (kc_kennylau):

that's an identity: \(\sin(x+y)=\sin x\cos y+\cos x\sin y\)

OpenStudy (kc_kennylau):

you just have to use this identity

OpenStudy (kc_kennylau):

you are finding \(\sin(x+y)\), which is equivalent to \(\sin x\cos y+\cos x\sin y\), so you substitute the numbers in there to find its value

OpenStudy (anonymous):

soooo, -4/5*15/7+15/5*-4/7?? :)

OpenStudy (kc_kennylau):

no...

OpenStudy (kc_kennylau):

ok draw two right triangles, one for x and one for y

OpenStudy (kc_kennylau):

you'll have to use PythThm

OpenStudy (anonymous):



OpenStudy (kc_kennylau):



OpenStudy (kc_kennylau):

find the other two sides

OpenStudy (anonymous):

okay, 3 and 16? but how is a side of a triangle negative?

OpenStudy (kc_kennylau):

because it's in the fourth quadrant and everything in the fourth quadrant is cool (8^2+15^2=17^2)

OpenStudy (anonymous):

lol so its negative because its cool? that makes sense :P lmao buttt was i right baout the 3 and 16? and what is that equation u put there??

OpenStudy (kc_kennylau):

it's not 16 it's 8 (8^2+15^2=17^2)

OpenStudy (kc_kennylau):

and can i finish my lunch first i'm so sorry

OpenStudy (anonymous):

lol yeah im in no rush, thanks for all of this. lol and lunch? its 11 o clock at night hereee ahaa

OpenStudy (kc_kennylau):

hi im back :)

OpenStudy (anonymous):

hey sorry i wasnt at my computer, but i finally found the answer to that one on google. because no matter how long i looked at all of the above and tried to figure it out, I COULDNT DO IT,, so i gave up and cheated :l

OpenStudy (kc_kennylau):

@kittiwitti1 maybe you can help him :D